5. Find the least squares line associated with the following data. Show all work for Find the l. any credit 5 71
Answer #1
x y xy x2
2.00 2.50 5.00 4.00
3.00 3.60 10.80 9.00
5.00 4.10 20.50 25.00
7.00 5.00 35.00 49.00
\sum 17.00 15.20 71.30 87.00

Η Σ 17/4 = 4.25

Η Σ9 – 15.2/4 = 3.80

Slope of Least squares line is given by

хи — rСо f.ru — fх 7 – 4

b =- 71.30 – 4* 4.25 * 3.80 -= 0.454237 87,00 – 4* 4.252

a = y - 17 = 3.80 – 0.454237 * 4.25 = 1.869493

Least square line is given by

y = a + br = 1.869493+ 0.4542370